OpenStudy (anonymous):

At a concession stand , 5 hot dogs and 2 hamburgers cost 11.25. find the cost of 1 hot dog and 1 hamburger

OpenStudy (anonymous):

im still having trouble trying to solve help?

OpenStudy (anonymous):

can u explain step by step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

asker, are you sure the question is complete?

OpenStudy (anonymous):

yes its complete

OpenStudy (anonymous):

oh no im sorry its not!! lol sorry here it goes i left out a part

OpenStudy (anonymous):

At a concession stand 5 hot dogs and 2 hamburgers cost 9.75, 2 hot dogs and 5 hamburgers cost 11.25 Find the cost of 1 hot dog and 1 hamburger.

OpenStudy (anonymous):

sorry for the mistake guys.

OpenStudy (anonymous):

let : d=cost of 1 hotdog h=cost of 1 hamburger so now you can write down your two equations...

OpenStudy (anonymous):

d+h=9.75 h+d=11.25?

OpenStudy (anonymous):

no... that's 5(hotdogs) + 2(hamburgers) = 9.75 or 5d + 2h = 9.75

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so what's the other equation involving 11.25 ???

OpenStudy (anonymous):

2h+5d=11.25

OpenStudy (anonymous):

.... and then?

OpenStudy (anonymous):

yes.... so your two equations are: 5d + 2h = 9.75 2d + 5h = 11.25

OpenStudy (anonymous):

wait.... the second equation is 2(hotdogs) + 5(hamburgers) = 11.25, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... do you know elimination or substitution to solve systems of equations?

OpenStudy (anonymous):

some what not quite, can you show it to me

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Take it like this... \[5d+2h=9.75------1.\]\[2d+5h=11.25----2.\]Now multiply 1 by 2 to get\[10d+4h=19.5-----3.\]Now multiply 2 by 5.\[10d+15h=56.25------4.\]Then minus 4 by 3..

OpenStudy (anonymous):

@Miyuru shows elimination of the "d" variable.. i'll let him continue.... :)

OpenStudy (anonymous):

i think its -5

OpenStudy (anonymous):

that u multiply by right isnt it -56.26

OpenStudy (anonymous):

im stuck. please help

OpenStudy (anonymous):

No I multiplied both sides by 5.. :)

OpenStudy (anonymous):

ok... continuing what was done, subtracting the two equations 10d + 4 h = 19.75 10d + 25h = 56.25 Let's take (equation 2) - (equation 1) you'd get: 21h = 36.5 you see how i got that?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

I have done a multiplication error in equation no 4 :D

OpenStudy (anonymous):

SO THEN YOU DIVIDE?

OpenStudy (anonymous):

NO UR GOOD MIYURU

OpenStudy (anonymous):

Thanks.. :D

OpenStudy (anonymous):

WHAT DID U GET FOR THE COST OF EACH MIYURU?

OpenStudy (anonymous):

I got h=1.7 and d=0.6

OpenStudy (anonymous):

OK, AWESOME. THANK YOU.

OpenStudy (anonymous):

mistake in the first equation: 10d + 4 h = 19.50 10d + 25h = 56.25 Let's take (equation 2) - (equation 1) you'd get: 21h = 36.75

OpenStudy (anonymous):

from here, 36.75/21 = 1.75, the cost of 1 hamburger. plug this cost into one of the original equations to get the cost of the hotdog.

OpenStudy (anonymous):

@spimentel dpalnc has corrected a mistake... :D